### v3

# Static and Dynamic BART for Rank-Order Data

Abstract
Ranking lists are often provided at regular time intervals in a range of applications, including economics, sports, marketing, and politics. Most popular methods for rank-order data postulate a linear specification for the latent scores, which determine the observed ranks, and ignore the temporal dependence of the ranking lists. To address these issues, novel nonparametric static (ROBART) and autoregressive (ARROBART) models are developed, with latent scores defined as nonlinear Bayesian additive regression tree functions of covariates.
To make inferences in the dynamic ARROBART model, closed-form filtering, predictive, and smoothing distributions for the latent time-varying scores are derived. These results are applied in a Gibbs sampler with data augmentation for posterior inference.
The proposed methods are shown to outperform existing competitors in simulation studies, static data applications to electoral data, stated preferences for sushi and movies, and dynamic data applications to economic complexity rankings of countries and weekly pollster rankings of NCAA football teams.
Keywords: Autoregressive Process; BART; Filtering and Smoothing; Rank-Order Data; Thurstone model
1 Introduction
Rank-order data emerge when an expert is asked to rank a finite set of $N$ items according to a certain criterion, such as personal preference or a predetermined statistic. The study of rank-order data is becoming popular in several fields, including political science (Gormley and Murphy, 2008), music preferences (Bradlow and Fader, 2001), economics (Fok et al., 2012), and sports analytics (Graves et al., 2003). In practice, the rankers are likely heterogeneous in their qualities or opinions, and it is often of interest to investigate how the rankings may be dependent on covariates of the ranked items (e.g., summary statistics). For example, for sports leagues such as the NCAA, NFL, or NBA, many experts provide ranking lists of teams by relying on their experiences and team-specific statistics, like the win percentage and the margin of victory. Moreover, reflecting the inherently dynamic nature of competitors’ abilities and human preferences, the experts’ rankings may change over time (e.g., Graves et al., 2003). Analogous considerations hold for rank-order data from other fields, such as marketing, politics, and economics.
In this article, we focus on models for rank-order data (Alvo and Philip, 2014; Liu et al., 2019). Specifically, we consider the family of Thurstone models, which assumes the existence of an unobserved evaluation score for each of the $N$ items whose noisy realizations determine the rankings provided by the $M$ experts. However, most of the existing models in this class are designed to study cross-sectional data and are inherently static, thus ignoring any temporal dependence.
Refer to caption Refer to caption
Figure 1: Left panel: Economic Complexity Index rankings for a selection of countries by year. Right panel: observed relative ranks for $N=7$ teams of the 2022 NCAA Division I Football season by AP Pollster Brian Howell.
Figure 1 shows the time-varying rankings of the Economic Complexity Index (ECI) (Simoes and Hidalgo, 2011) for 8 countries from 1965 to 2015, and the 2022 weekly ranking of NCAA teams by a single pollster. The changes in rankings over time are heterogeneous in magnitude, with both large and small shifts, as well as a few series having a stable ranking over time. Overall, these features highlight both slow evolution and large breaks due to events such as global crises or small events at a country level. The NCAA dataset shows similar features, with apparently smoother dynamics due to the smaller number of items. Section 5 includes an application to these datasets among the others.
The main challenge in investigating and forecasting this class of data is the constraint among the items in a ranking list, which prevents the use of equation-by-equation modeling and calls for specific statistical tools. However, the existing rank-order models fall short of the flexibility necessary (i) to address the highly nonlinear relationship between the item’s features and observed rankings and (ii) to forecast time series characterized by possibly large changes.
This article contributes to the literature on rank-order models by proposing two new methods for addressing these issues. Specifically, we leverage the Bayesian additive regression trees (BART, see Chipman et al., 2010) framework to define a flexible nonparametric static model for rank-order data, named rank-order Bayesian additive regression trees (ROBART). This model is an extension of the linear rank-order model of Li et al. (2022) to a setting with non-linearities and covariates that are specific to both items and rankers. Then, we extend this framework to a dynamic model called autoregressive rank-order Bayesian additive regression trees (ARROBART), which accounts for the persistence in time series of rank-order data (e.g., see Fig. 1).
The most widely used families of models for rank data are the Thurstone (Thurstone, 1927) and Plackett-Luce (Luce, 1959; Plackett, 1975) models, which differ in the distribution of the noise in the random utility representations (normal and Gumbel, respectively) and in the ordering interpretation (joint or sequential). Relying on normal regression models makes the Thurstone family easier to interpret and extend in several directions. However, the estimation of these models is generally difficult, especially for large numbers of ranked items, due to the high-dimensional integral form of the likelihood function. Within the Thurstone family, the Bayesian approach coupled with data augmentation methods provides an advantageous computational alternative that also allows for the introduction of covariates (Yu, 2000). Johnson et al. (2002) developed a hierarchical Bayesian Thurstone model for rank-order data from known groups of evaluators, whereas Li et al. (2022) introduced ranker-specific error variances and a Dirichlet process mixture prior for ranker-specific effects of covariates. Recently, Gu and Yu (2023) introduced a Bayesian spatial autoregressive model to learn ranking data in social networks and to account for social dependencies among individuals. We contribute to this literature by proposing two new Thurstone models, a static and a dynamic one, where a Bayesian nonparametric approach is used to capture the possibly nonlinear impact of covariates. Moreover, we extend the proposed methods to address incomplete data (i.e., partial rankings) in the form of top-$q$ rankings, which frequently happens when rankers assess many items.
The study of time series of rankings is less developed but has attracted increasing attention in recent years. Most existing dynamic ranking models are generalizations of the Plackett-Luce model (Bradlow and Fader, 2001; Grewal et al., 2008; Glickman and Hennessy, 2015; Henderson and Kirrane, 2018), whereas this article provides a dynamic extension to the Thurstone family. We propose a novel time-varying nonparametric approach that considers an autoregressive process for the latent scores and uses a specification based on regression trees to flexibly model the impact of the lagged response and the external covariates. Its linear counterpart is also introduced as a benchmark.
The possibility of having large deviations as well as small changes in the time series of ranking lists, as shown in Fig. 1, calls for the use of flexible tools to model the dynamics of the latent “scores” driving the ranking of each item. To address these issues, we introduce the BART specification to model the conditional mean of latent scores. We leverage the good performance of this approach for fitting and forecasting regression models (e.g., see Hill et al., 2020) in order-statistics models to design static and dynamic nonparametric frameworks for rank-order data. Despite its advantages, BART has rarely been used in time series or state space models. Recent advances include Bayesian additive vector autoregressive trees for multiple time series (Huber and Rossini, 2022), and an extension to a mixed-frequency context for nowcasting in the presence of extreme observations (Huber et al., 2023) . Our contribution to this literature concerns the derivation of new closed-form filtering, predictive, and smoothing distributions for the latent time-varying scores in the ARROBART model. Furthermore, we show the proportionality of these distributions to finite mixtures.
The proposed ROBART and ARROBART models are tested on static and dynamic synthetic datasets, respectively, highlighting the superior performance against their linear counterparts as well as other existing methods. The proposed static models are then applied to three datasets about sushi and movie preference data and election results, followed by two applications of the dynamic model to the rankings from the Economic Complexity Index (ECI) rankings and the 2022 NCAA pollsters displayed in Fig. 1. We find evidence supporting the superior performance of ROBART and ARROBART in the static and dynamic settings, respectively, relative to their linear counterparts, which suggests that it is important to account for both temporal persistence and nonlinearities.
The remainder of the article is organized as follows: Section 2 provides an overview of the Thurstone model and illustrates the ROBART and ARROBART models. It also provides a theorem describing the filtering and smoothing distributions for ARROBART. Section 3 presents a Bayesian approach to inference. Section 4 investigates the performance of our methods on simulated data. Section 5 includes applications to three static ranking data sets and a time series of Economic Complexity Index rankings. Section 6 concludes the article.
2 Nonparametric Thurstone Models for Rank-Order Data
This section introduces notation, then presents a new class of nonparametric models for rank-order data that extends the Thurstone model (Thurstone, 1927) by allowing for flexible, possibly nonlinear functions of the covariates. Then, we extend the framework to account for temporal dependence and persistence for investigating and forecasting time series data.
A full ranking of $N$ items is a map $\tau:\mathds{A}\to\mathcal{P}_{N}$ from a finite, ordered set of items, $\mathds{A}=\{a_{1},\ldots,a_{N}\}$ to the space of $N$-dimensional permutations, $\mathcal{P}_{N}$. Let ##\## denote the cardinality of a set. With a slight abuse of notation, we denote a full ranking of $N$ items by the vector $\tau=(\tau_{1},\ldots,\tau_{N})^{\prime}\in\mathcal{P}_{N}$, where $\tau_{i}=\tau(i)$ is the rank of item $a_{i}\in\mathds{A}$, for $i=1,\ldots,N$, and $\tau_{i}<\tau_{j}$ means that item $a_{i}$ is preferred to $a_{j}$, written $a_{i}\succ a_{j}$ since it is assigned a lower rank. Rankings and pairwise comparisons are related, as, given a full ranking $\tau$, it holds $[a_{i}\succ a_{j}]\iff\tau_{i}<\tau_{j}$ for each $i,j\in\{1,\ldots,N\}$. Therefore, a full ranking $\tau$ is an ordered $N$-tuple representing a set of non-contradictory pairwise relationships in $\mathds{A}$, and as such, it encodes an ordered preference for all the items in $\mathds{A}$.
In this article, we consider a setting where $N$ items are ordered by $M$ rankers. For each item $i=1,\ldots,N$ and ranker $j=1,\ldots,M$, the $K_{x}$-dimensional vector of covariates is denoted by $\mathbf{X}_{ij}\in\mathds{R}^{K_{x}}$. To account for external information that might affect the rankers’ decision process, it is possible to include item- and ranker-specific covariates, $\mathbf{x}_{i}^{a}\in\mathds{R}^{K_{a}}$ and $\mathbf{x}_{j}^{r}\in\mathds{R}^{K_{r}}$, as well as covariates specific to both items and rankers, $\mathbf{w}_{ij}\in\mathds{R}^{K_{w}}$. In the dynamic case, this setup is replicated for each of the $T$ time periods. Finally, for any vector $\mathbf{z}=(z_{1},\ldots,z_{N})^{\prime}\in\mathds{R}^{N}$, we use $\operatorname{rank}(\mathbf{z})=[z_{i_{1}}\succ\ldots\succ z_{i_{N}}]$ to denote the full ranking of $z_{i}$’s in decreasing order of preference, that is, $z_{i_{1}}<\ldots<z_{i_{N}}$.
2.1 Overview of Bayesian Additive Regression Trees
A Bayesian additive regression tree (BART) specification for a function $f:\mathds{R}^{K_{x}}\to\mathds{R}$ is an approximation of $f$ by means of a sum of regression trees. Let us denote by $\mathcal{T}$ a binary tree consisting of a set of interior node decision rules and a set of terminal nodes (or leaves), and let $\bm{\mu}=\{\mu_{1},\ldots,\mu_{b}\}$ be a set of parameter values associated with each of the $b$ terminal nodes of $\mathcal{T}$. A regression tree takes covariate values as input and produces a fitted outcome value given by the parameter associated with a terminal node of the tree. This outcome is obtained by dividing the domain of the covariate $\mathbf{x}_{i}=(x_{i1},\ldots,x_{iK_{x}})^{\prime}\in\mathds{R}^{K_{x}}$ into disjoint regions $\{B_{\ell}\}_{\ell=1}^{b}$ using a sequence of binary rules of the form $\{\mathbf{x}_{i}\in A_{h}\}$ versus $\{\mathbf{x}_{i}\notin A_{h}\}$, where $h\in\mathcal{I}$ indexes the internal nodes of the tree. Most commonly, $A_{h}$ is defined by a splitting variable $k_{h}$ and splitting point $c_{h}$, such as $A_{h}=\{\mathbf{x}_{i}:x_{ik_{h}}\leq c_{h}\}$. Then, $\mathbf{x}_{i}$ is assigned the value $\mu_{\ell}$, with $\ell\in\{1,\ldots,b\}$, associated with a single terminal node $B_{\ell}$ of $\mathcal{T}$ defined by a sequence of decision rules.
A tree $\mathcal{T}$ and a collection of leaf parameters $\bm{\mu}$ define a step function $g(\mathbf{x}_{i}|\mathcal{T},\bm{\mu})$ that assigns value $\mu_{\ell}\in\bm{\mu}$ to $\mathbf{x}_{i}$, thus resulting in:
$g(\mathbf{x}_{i}|\mathcal{T},\bm{\mu})=\sum_{\ell=1}^{b}\mu_{\ell}\mathbb{I}(\mathbf{x}_{i}\in B_{\ell}),$
where $\mathbb{I}(\cdot)$ is the indicator function and $\{B_{\ell}\}_{\ell=1}^{b}$ is the partition of the domain of the covariate $\mathbf{x}_{i}$ associated to the tree $\mathcal{T}$. The BART specification consists in approximating the function $f$ through a sum of $S$ trees, as follows:
$f(\mathbf{x}_{i})\approx\sum_{s=1}^{S}g_{s}(\mathbf{x}_{i}|\mathcal{T}_{s},\bm{\mu}_{s})=\sum_{s=1}^{S}\sum_{\ell=1}^{b_{s}}\mu_{\ell,s}\mathbb{I}(\mathbf{x}_{i}\in B_{\ell,s}),$ (1)
where $g_{s}$ is a tree function that assigns $\mu_{\ell,s}\in\bm{\mu}_{s}$ to $\mathbf{x}_{i}$ for each binary regression tree $\mathcal{T}_{s}$.
2.2 Static Model: Rank-Order BART
The Thurstone model can be described as a random utility model (e.g., see Walker and Ben-Akiva, 2002), where the full ranking $\tau$ is driven by a latent score, $\mathbf{z}$, such that for any two items $i_{1},i_{2}$, it holds $z_{i_{1}}<z_{i_{2}}$ if and only if $i_{1}\succ i_{2}$. The latent vector $\mathbf{z}$ is assumed to follow a multivariate normal distribution with mean $\bm{\gamma}=(\gamma_{1},\ldots,\gamma_{N})^{\prime}$ representing the score of each item:
$\tau=\operatorname{rank}(\mathbf{z}),\qquad z_{i}=\gamma_{i}+\varepsilon_{i},\qquad\varepsilon_{i}\sim\mathcal{N}(0,\sigma_{\varepsilon}^{2}).$ (2)
[…]
Consider a framework where no temporal observations are available or the time dependence is irrelevant to the analysis. Moreover, we assume each individual in a collection of $M$ professionals or rankers provides a ranking list of the same set of $N$ items, as is common in many real-world situations such as sports analytics and customer reviews. To investigate data without temporal dependence, we define a new static model that extends the Thurston approach in Eq. (2) by including the BART specification from Eq. (1) for the conditional mean of the latent scores. Specifically, we generalise Eq. (2) by assuming ranker-specific scores, $\mathbf{z}_{j}$, and, for each pair $(i,j)$, we replace the intercept $\gamma_{ij}$ with an unknown function of the covariates, $f$. Moreover, to cope with potential nonlinearities, interactions, and discontinuities in the rankers’ dynamic scores, we adopt a nonparametric approach based on a BART specification for the latent regression function, $f$.
For each item $i=1,\ldots,N$ and ranker $j=1,\ldots,M$, let $\mathbf{X}_{ij}=(\mathbf{x}_{i}^{a\prime},\mathbf{x}_{j}^{r\prime},\mathbf{w}_{ij}^{\prime})^{\prime}\in\mathds{R}^{K_{x}}$ be a covariate vector. The Rank-Order BART model (ROBART) is defined as: [^2] Assuming that the latent scores are item-ranker specific implicitly means assuming each ranker may make a different assessment of the items. Instead, when a single score is postulated for each item, it would imply postulating that each ranker makes a (noisy) evaluation of a unique and objective ranking of the items. Moreover, the former approach allows us to include ranker-specific covariates, $\mathbf{x}_{j}^{r}$, and object-ranker-specific covariates, $\mathbf{w}_{ij}$, as factors possibly affecting the latent scores.
$\begin{split}\tau_{j}=\operatorname{rank}(\mathbf{z}_{j}),\qquad z_{ij}&=f(\mathbf{X}_{ij})+\varepsilon_{ij},\qquad\varepsilon_{ij}\mathbin{\overset{iid}{\kern 0.0pt\sim}}\mathcal{N}(0,1),\\ f(\mathbf{X}_{ij})&=\sum_{s=1}^{S}\sum_{\ell=1}^{b_{s}}\mu_{\ell,s}\mathbb{I}(\mathbf{X}_{ij}\in B_{\ell,s}),\end{split}$ (3)
[…]
$\begin{split}L(\tau_{1},\ldots,\tau_{M}|\bm{\theta})&=\prod_{j=1}^{M}\int_{\mathds{R}^{N}}(2\pi)^{-\frac{N}{2}}\exp\Bigg{(}-\frac{\big{\|}\mathbf{z}_{j}-f(\mathbf{X}_{ij})\big{\|}^{2}}{2}\Bigg{)}\delta_{\operatorname{rank}(\mathbf{z}_{j})}(\tau_{j})\>\mathrm{d}\mathbf{z}_{j}.\end{split}$ (4)
2.3 Dynamic Model: AutoRegressive Rank-Order BART
In many real-world cases, multiple professionals continuously evaluate the items and provide their ranking lists periodically over time. The latent scores that a ranker assigns to each item, which ultimately determine the professional’s ranking lists observed by the researcher, evolve over time to reflect the change in the information set and in the features of the items. Nonetheless, most of the existing models for rank-order data disregard the temporal dimension of the data, thus missing information relating to the serial correlation in the rankers’ latent scores and observed ranks. We aim to address this issue by proposing a modification of the Thurstone model that accounts for the temporal persistence of the ranking lists as implied by an evolution of the underlying rankers’ scores.
Suppose that each ranker provides a ranking of items at each time $t=1,\ldots,T$, and let $\tau_{j,t}$ denote the observed ranking list of ranker $j$ in period $t$. Moreover, let $\mathbf{z}_{j,t}=(z_{1j,t},\ldots,z_{Nj,t})^{\prime}$, where $z_{ij,t}$ denotes the unobserved ranker $j$’s score for item $i$, such that $z_{i_{1}j,t}>z_{i_{2}j,t}$ if and only if $i_{1}\prec i_{2}$ for ranker $j$ at time $t$. In this setting, each latent score $z_{ij,t}$ can be modeled as an unknown function $f$ of its first lag and the external covariates. This is done by considering the covariate vector $\mathbf{X}_{ij,t}=(z_{ij,t-1},\mathbf{x}_{i,t}^{a\prime},\mathbf{x}_{j,t}^{r\prime},\mathbf{w}_{ij,t}^{\prime})^{\prime}\in\mathds{R}^{K_{x}}$, where $\mathbf{x}_{i,t}^{a}$, $\mathbf{x}_{j,t}^{r}$, and $\mathbf{w}_{ij,t}$ are the item-, ranker-, and item/ranker-specific variables.
Therefore, in the presence of time series rank-order data where temporal persistence is a relevant feature, we extend the ROBART model in Eq. (3) to a dynamic framework by defining an AutoRegressive Rank-Order BART model with exogenous covariates (ARROBARTX) as follows:
$\begin{split}\tau_{j,t}=\operatorname{rank}(\mathbf{z}_{j,t}),\qquad z_{ij,t}&=f(\mathbf{X}_{ij,t})+\varepsilon_{ij,t},\qquad\varepsilon_{ij,t}\mathbin{\overset{iid}{\kern 0.0pt\sim}}\mathcal{N}(0,1),\\ f(\mathbf{X}_{ij,t})&=\sum_{s=1}^{S}\sum_{\ell=1}^{b_{s}}\mu_{\ell,s}\mathbb{I}(\mathbf{X}_{ij,t}\in B_{\ell,s}),\end{split}$ (5)
where the regression function $f$ is approximated by a BART specification. A special case of the ARROBARTX in Eq. (5) is an AutoRegressive Rank-Order BART model without covariates (ARROBART) and with one lag, which is obtained by assuming $\mathbf{X}_{ij,t}=\mathbf{z}_{j,t-1}$. The ARROBART model can be interpreted as a hidden Markov model (HMM, see Cappé et al., 2005) with measurement and transition density given by, respectively:
$\tau_{j,t}|\mathbf{z}_{j,t}\sim\delta_{\operatorname{rank}(\mathbf{z}_{j,t})}(\tau_{j,t}),\text{ and }\mathbf{z}_{j,t}|\mathbf{z}_{j,t-1}\sim\mathcal{N}_{N}(\mathbf{z}_{j,t}|f(\mathbf{z}_{j,t-1}),I_{N}),$
where $f$ is approximated by a BART and $I_{N}$ denotes the $N$-dimensional identity matrix. Let $\bm{\tau}=\{\tau_{j,t}:j\in\{1,\dots,M\},t\in\{1,\dots,T\}\}$ denote the collection of all the observations and $\bm{\theta}$ be the collection of all parameters. The ARROBART likelihood function is:
$L(\bm{\tau}|\bm{\theta})=\prod_{j=1}^{M}\underbrace{\int_{\mathds{R}^{N}}\dots\int_{\mathds{R}^{N}}}_{\text{T+1 integrals}}\left(\prod_{t=1}^{T}p(\tau_{j,t}|\mathbf{z}_{j,t})p(\mathbf{z}_{j,t}|\mathbf{z}_{j,t-1},\bm{\theta})\right)p(\mathbf{z}_{j,0}|\bm{\theta})\>\mathrm{d}\mathbf{z}_{j,0}\dots\mathrm{d}\mathbf{z}_{j,T},$ (6)
where $\mathbf{z}_{j,0}$ is the initial value of the latent vector. As the ARROBART model postulates independence across rankers $j$, hereafter, without loss of generality, we focus on the case $M=1$ and drop the corresponding index.
The temporal dependence among the latent variables introduces a filtering problem in the estimation of the ARROBART model, as making inferences on the latent vector $\mathbf{z}_{t}$ requires the derivation of appropriate filtering and smoothing distributions.
Let us first introduce some additional notation. For a variable $x\in\mathcal{D}\subseteq\mathds{R}$, the BART specification induces a partition $\{C_{k}\}_{k=1}^{K}$ of the domain $\mathcal{D}$, such that the (approximated) function has constant value $\tilde{\mu}_{k}$ over each region $C_{k}$. We refer to the Supplement for proof and give a sketch below. First, notice that, for each tree $s\in\{1,\ldots,S\}$ defined in eq. (1), the collection of intervals $\{B_{\ell,s}\}_{\ell=1}^{b_{s}}$ is a partition of $\mathcal{D}$ and the $\ell$th leaf of the $s$th tree is associated to the scalar coefficient $\mu_{\ell,s}\in\mathds{R}$. Moreover, we can associate each tuple (j1,,jS)
s=1S
{1,,bs}
[…]
one-to-one to an index $k\in\{1,\ldots,K\}$, where $K=\prod_{s=1}^{S}b_{s}$ is the total number of the tuples. Then, for each tuple $(j_{1},\ldots,j_{S})$ and associated $k$ index, define $C_{k}=B_{j_{1},1}\cap\ldots\cap B_{j_{S},S}$ and let $\tilde{\mu}_{k}=\sum_{s=1}^{S}\mu_{j_{s},s}$ be the associated coefficient. Note that the intervals $C_{k}$ can be empty ($C_{k}=\emptyset$) or unbounded, that is $C_{k}=(-\infty,a)$ or $C_{k}=(b,\infty)$ for some $a,b\in\mathcal{D}$.
For simplicity, we introduce some shorthand notation. For a vector $\tau_{i}$ and a set $A_{i}$, $i=1,\ldots,t$, we define the collections $\tau_{1:t}=\{\tau_{1},\ldots,\tau_{t}\}$ and $A_{1:t}=\{A_{1},\ldots,A_{t}\}$. Let $\mathbf{k}=(k_{1},\ldots,k_{N})^{\prime}\in\{1,\ldots,K\}^{N}$ denote a tuple or multi-index and $\bm{\tilde{\mu}}_{\mathbf{k}}=(\tilde{\mu}_{k_{1}},\ldots,\tilde{\mu}_{k_{N}})^{\prime}$ be the vector of leaf parameters indexed by $\mathbf{k}$. Then, a summation over multiple indices is written as
$\sum_{\mathbf{k}}a_{\mathbf{k}}=\sum_{\mathbf{k}\in\{1,\ldots,K\}^{N}}a_{\mathbf{k}}=\sum_{k_{1}=1}^{K}\dots\sum_{k_{N}=1}^{K}a_{k_{1},\ldots,k_{N}}.$
We define $A_{t}$ as the set of vectors $\mathbf{z}_{t}$ whose elements follow the same ordering as in the raking list $\tau_{t}$, that is:
$A_{t}=\big{\{}\mathbf{z}_{t}\in\mathds{R}^{N}:z_{j,t}<z_{i,t}\iff\tau_{j,t}<\tau_{i,t},\;\;\forall\,i,j=1,\ldots,N\big{\}}.$
[…]
i=1N
Cki
[…]
and $C_{\mathbf{k},t}=C_{\mathbf{k}}\cap A_{t}$, for $t=1,\ldots,T$. Therefore, for any $\mathbf{z}_{t}\in C_{\mathbf{k},t}$, all the element-wise constraints $\{C_{k_{i}}\}_{i=1}^{N}$ and the ordering restriction $A_{t}$ are satisfied.
Theorem 1 below derives the filtering, predictive, and smoothing distributions for the latent vector $\mathbf{z}_{j,t}$, then Corollary 1 shows that they are proportional to finite mixtures with time-varying weights and components. We refer to the Supplement for the proofs.
Theorem 1.
The filtering distribution is given by
$p(\mathbf{z}_{t}|\tau_{1:t})\propto\mathbb{I}(\mathbf{z}_{t}\in A_{t})\sum_{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{k}\in\{1,\ldots,K\}^{N}}\mathcal{N}_{N}(\mathbf{z}_{t}|\bm{\tilde{\mu}}_{\mathbf{k}},I_{N})q_{\mathbf{k},t}(A_{1:t-1}),$ (7)
where $A_{t}$ is the set of possible vectors $\mathbf{z}_{t}$ that agree with the rank-order given by $\tau_{t}$, $\bm{\tilde{\mu}}_{\mathbf{k}}$ are mean vectors corresponding to possible combinations of regions $C_{k_{1}},\ldots,C_{k_{N}}$ in which the latent variable $\mathbf{z}_{t-1}$ can be located, and for $t>1$,
$q_{\mathbf{k},t}(A_{1:t-1})$ $=\sum_{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{m}\in\{1,\ldots,K\}^{N}}q_{\mathbf{m},t-1}(A_{1:t-2})\int_{\mathbf{C}_{\mathbf{k},t-1}}\mathcal{N}_{N}(\mathbf{z}_{t-1}|\bm{\tilde{\mu}}_{\mathbf{m}},I_{N})\>\mathrm{d}\mathbf{z}_{t-1},$
whereas, at time $t=1$ there is no dependence on $A_{t}$ and we have:
$q_{\mathbf{k},1}$ $=\int_{\mathbf{C}_{\mathbf{k}}}\mathcal{N}_{N}(\underline{\mathbf{z}}_{0}|\underline{\mathbf{z}}_{prior},I_{N})\>\mathrm{d}\underline{\mathbf{z}}_{0}.$
The one-step-ahead predictive distribution, for each $t=1,\ldots,T-1$, is
$p(\mathbf{z}_{t+1}|\tau_{1:t})\propto\sum_{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{k}\in\{1,\ldots,K\}^{N}}\mathcal{N}_{N}(\mathbf{z}_{t+1}|\bm{\tilde{\mu}}_{\mathbf{k}},I_{N})q_{\mathbf{k},t+1}(A_{1:t}).$ (8)
The smoothing distribution is $p(\mathbf{z}_{t}|\tau_{1:T})\propto p(\mathbf{z}_{t}|\tau_{1:t})r_{t}(\mathbf{z}_{t})$, where $r_{t}(\mathbf{z}_{t})$ embeds all information from the periods $t^{\prime}=t+1,\ldots,T$ and is defined recursively for each $t<T$ as
$r_{t}(\mathbf{z}_{t})$ $=\int_{A_{t+1}}r_{t+1}(\mathbf{z}_{t+1})\mathcal{N}_{N}\Big{(}\mathbf{z}_{t+1}\big{|}\sum_{s=1}^{S}g_{s}(\mathbf{z}_{t}),I_{N}\Big{)}\>\mathrm{d}\mathbf{z}_{t+1}$
$=\int_{A_{t+1}}r_{t+1}(\mathbf{z}_{t+1})\mathcal{N}_{N}\Big{(}\mathbf{z}_{t+1}\big{|}\sum_{\mathbf{k}\in\{1,\ldots,K\}^{N}}\bm{\tilde{\mu}}_{\mathbf{k}}\mathbb{I}(\mathbf{z}_{t}\in C_{\mathbf{k}}),I_{N}\Big{)}\>\mathrm{d}\mathbf{z}_{t+1},$
where tree function $g_{s}$ is defined as $g_{s}(\mathbf{x}|\mathcal{T}_{s},\bm{\mu}_{s})=\sum_{\ell=1}^{b_{s}}\mu_{\ell,s}\mathbb{I}(\mathbf{x}\in B_{\ell,s})$ and it holds $\sum_{s=1}^{S}g_{s}(\mathbf{x}|\mathcal{T}_{s},\bm{\mu}_{s})=\sum_{\mathbf{k}\in\{1,\ldots,K\}^{N}}\bm{\tilde{\mu}}_{\mathbf{k}}\mathbb{I}(\mathbf{x}\in C_{\mathbf{k}})$. At the final time $t=T$, it is given by $r_{T}(\mathbf{z}_{T})=1$.
Corollary 1 shows that both the filtering and smoothing distributions are proportional to finite mixtures. These alternative representations for the filtering and smoothing distributions can be used to design a more efficient algorithm for posterior inference based on a data augmentation approach coupled with a Metropolis-Hastings step (see Section 3).
Corollary 1 (Mixture representations).
At each time $t=1,\ldots,T$, the filtering distribution is proportional to the finite mixture:
$p(\mathbf{z}_{t}|\tau_{1:t})$ $\propto\sum_{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{k}\in\{1,\ldots,K\}^{N}}w^{F}_{\mathbf{k},t}u^{F}_{\mathbf{k},t}(\mathbf{z}_{t}),$ (9)
with weights $w^{F}_{\mathbf{k},t}$ and components $u^{F}_{\mathbf{k},t}(\mathbf{z}_{t})=\bar{u}^{F}_{\mathbf{k},t}(\mathbf{z}_{t})/n_{\mathbf{k},t}$, where
$\bar{u}^{F}_{\mathbf{k},t}(\mathbf{z}_{t})$ $=\mathbb{I}(\mathbf{z}_{t}\in A_{t})\mathcal{N}_{N}(\mathbf{z}_{t}|\bm{\tilde{\mu}}_{\mathbf{k}},I_{N}),$
$n_{\mathbf{k},t}$ $=\int_{\mathds{R}^{N}}\bar{u}^{F}_{\mathbf{k},t}(\mathbf{z}_{t})\>\mathrm{d}\mathbf{z}_{t}=\int_{A_{t}}\mathcal{N}_{N}(\mathbf{z}_{t}|\bm{\tilde{\mu}}_{\mathbf{k}},I_{N})\>\mathrm{d}\mathbf{z}_{t},$
$w^{F}_{\mathbf{k},t}$ $=\frac{q_{\mathbf{k},t}(A_{1:t-1})n_{\mathbf{k},t}}{\sum_{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{m}\in\{1,\ldots,K\}^{N}}q_{\mathbf{m},t}(A_{1:t-1})n_{\mathbf{m},t}}.$
At each time $t=1,\ldots,T$, the smoothing distribution is proportional to the finite mixture:
$p(\mathbf{z}_{t}|\tau_{1:T})$ $\propto\sum_{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{k}\in\{1,\ldots,K\}^{N}}w^{S}_{\mathbf{k},t}u^{S}_{\mathbf{k},t}(\mathbf{z}_{t}),$ (10)
with weights $w^{S}_{\mathbf{k},t}$ and components $u^{S}_{\mathbf{k},t}(\mathbf{z}_{t})=\bar{u}^{S}_{\mathbf{k},t}(\mathbf{z}_{t})/m_{\mathbf{k},t}$, where
$\bar{u}^{S}_{\mathbf{k},t}(\mathbf{z}_{t})$ $=r_{t}(\mathbf{z}_{t})\mathbb{I}(\mathbf{z}_{t}\in A_{t})\mathcal{N}_{N}(\mathbf{z}_{t}|\bm{\tilde{\mu}}_{\mathbf{k}},I_{N}),$
$m_{\mathbf{k},t}$ $=\int_{\mathds{R}^{N}}\bar{u}^{S}_{\mathbf{k},t}(\mathbf{z}_{t})\>\mathrm{d}\mathbf{z}_{t},$
$w^{S}_{\mathbf{k},t}$ $=\frac{q_{\mathbf{k},t}(A_{1:t-1})m_{\mathbf{k},t}}{\sum_{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{\ell}\in\{1,\ldots,K\}^{N}}q_{\bm{\ell},t}(A_{1:t-1})m_{\bm{\ell},t}},\qquad\mathbf{k}\in\{1,\ldots,K\}^{N}.$
Notice that $\bar{u}^{F}_{\mathbf{k},t}(\mathbf{z}_{t})$ is a multivariate normal distribution restricted such that the elements satisfy the specific ordering encoded by $A_{t}$.
3 Bayesian Inference
3.1 Prior on the BART Parameters
The BART specification is completed by choosing a prior for the parameters of the sum-of-trees model, namely, the binary trees $\mathcal{T}_{s}$ and the terminal node parameters $\bm{\mu}_{s}$. We assume:
$p(\mathcal{T}_{1},\bm{\mu}_{1},\ldots,\mathcal{T}_{S},\bm{\mu}_{S})=\prod_{s=1}^{S}p(\bm{\mu}_{s}|\mathcal{T}_{s})p(\mathcal{T}_{s})=\prod_{s=1}^{S}\Big{[}\prod_{\ell=1}^{b_{s}}p(\mu_{\ell,s}|\mathcal{T}_{s})\Big{]}p(\mathcal{T}_{s}),$ (11)
where $\mu_{\ell,s}\in\bm{\mu}_{s}$. To deal with the potential overfitting of the nonparametric BART specification, we follow Chipman et al. (2010) and choose a regularization prior for the tree structure and terminal nodes that constrains the size and fit of each tree so that each one contributes only a small part to the overall fit. This prior strongly encourages each tree to be a weak learner, with leaf parameters close to zero.
The marginal prior for each tree structure, $p(\mathcal{T}_{s})$, is specified by three parts. First, the probability that a node at depth $d=0,1,\ldots$ is nonterminal, given by $\underline{\alpha}(1+d)^{-\underline{\beta}}$, for $\underline{\alpha}\in(0,1)$, $\underline{\beta}\in[0,\infty)$. Smaller (larger) values of $\underline{\alpha}$ ($\underline{\beta}$) impose a larger penalty on more complex tree structures. We use the default choice proposed by Chipman et al. (2010), that is $\underline{\alpha}=0.95$ and $\underline{\beta}=2$. Second, the distribution of the splitting variable at each interior node is assumed to be discrete uniform over all the covariates. Third, the distribution on the splitting rule assignment in each interior node, $c_{h}$, conditional on the splitting variable, is uniform on the discrete set of available splitting values.
The conditional prior on the terminal leaf parameters, $p(\mu_{\ell,s}|\mathcal{T}_{s})$, is assumed to be the conjugate normal distribution $\mathcal{N}(\mu_{\mu},\sigma^{2}_{\mu})$, where the hyperparameters are chosen such that $\mathcal{N}(S\mu_{\mu},S\sigma^{2}_{\mu})$ assigns substantial probability to the interval $(z_{min},z_{max})$. In line with Chipman et al. (2010), the prior on $\mu_{\ell,s}$ is constructed with $\mu_{\mu}=0$ and $\sigma_{\mu}=(z_{max}-z_{min})/(2k_{\mu}\sqrt{S})$, where $k_{\mu}$ is a hyperparameter set to 2, and $z_{max}$ and $z_{min}$ are the maximum and minimum of initial values of the latent outcome.
3.2 Posterior Sampling for ARROBART
To draw samples from the joint posterior distribution, a Markov Chain Monte Carlo (MCMC) algorithm is provided. The Gibbs sampler iterates over the following steps:
  1. 1)
    sample the path of the latent variables $\mathbf{z}=\{\mathbf{z}_{1},\ldots,\mathbf{z}_{T}\}$, recursively over time;
  2. 2)
    sample the tree and leaf parameters from $p(\mathcal{T}_{s},\bm{\mu}_{s}|\mathcal{T}_{-s},\bm{\mu}_{-s},\mathbf{z})$, for $s=1,\ldots,S$:
    1. a.
      compute the $NJT$-dimensional vector of partial residuals, $R_{s}$, as
      $R_{s,ijt}={z}_{ijt}-\sum_{\begin{subarray}{c}k=1\\ k\neq s\end{subarray}}^{S}g_{k}(\mathbf{X}_{ij,t}|\mathcal{T}_{k},\bm{\mu}_{k}),$
      and notice that $p(\mathcal{T}_{s},\bm{\mu}_{s}|\mathcal{T}_{-s},\bm{\mu}_{-s},\mathbf{z})=p(\mathcal{T}_{s},\bm{\mu}_{s}|R_{s})$;
    2. b.
      since $p(\mathcal{T}_{s}|R_{s})$ has a closed form, sample the trees structure from $p(\mathcal{T}_{s}|R_{s})$ using a Metropolis-Hastings algorithm;
    3. c.
      sample the terminal nodes from $p(\bm{\mu}_{s}|\mathcal{T}_{s},R_{s})$, which is a set of independent draws from a normal distribution.
Sampling the path of the latent variables using the filtering and smoothing recursions in Theorem 1 is highly computationally intensive. This is due to (i) evaluation of many Gaussian integrals over very small subspaces of $\mathds{R}^{N}$ and (ii) drawing from an $N$-dimensional Gaussian distribution truncated such that its elements satisfy a given ordering constraint. Both issues worsen as the number of items, $N$, increases. To overcome these computational issues, for each period $t=1,\ldots,T$, we approximate the joint posterior distribution of the vector $\mathbf{z}_{t}$ and derive the element-wise posterior full conditional distribution for each $i=1,\ldots,N$, conditioning on all the other $j\neq i$ elements. Therefore, we replace sampling from a joint distribution by sampling from a collection of full conditional distributions, which results in a full Gibbs sampler of the latent scores. This approach is similar to the standard technique used in rank-order data models (e.g., Hoff, 2009, Ch. 12). We refer to the Supplement for a detailed description of the sampling method.
Steps 2b and 2c concerning the static parameters of the tree structure are standard in the BART literature and follow directly from the algorithm in Chipman et al. (2010). The tree structure is updated by proposing one of four moves: growing a terminal node, pruning a pair of terminal nodes, changing a nonterminal rule, and swapping a rule between parent and child. Then, given the tree structure, the terminal node parameters are sampled directly from their normal conditional posterior distribution.
4 Simulation Studies
This section investigates the performance of the proposed methods on synthetic data. In particular, Section 4.1 compares the ROBART model to the methods for rank-order data introduced by Li et al. (2022). Then, Section 4.2 presents a comparative study between ARROBART and a range of models that account for (or ignore) the possible temporal persistence of the data.
4.1 Comparison of ROBART and Linear Models
4.1.1 Static Simulation Study with Full Rankings
The performance of the ROBART model defined in Section 2.2 is compared to the methods for rank-order data introduced by Li et al. (2022) by using the normalized Kendall tau distance (Kendall, 1938). Given two ranking lists $\hat{\tau}$ and $\tau$, the normalized Kendall tau distance is defined as the proportion of pairwise disagreements between them:
$K_{n}(\hat{\tau},\tau)=\frac{\#\big{\{}(i,j):i<j,\>[\hat{\tau}_{i}<\hat{\tau}_{j}\wedge\tau_{i}>\tau_{j}]\vee[\hat{\tau}_{i}>\hat{\tau}_{j}\wedge\tau_{i}<\tau_{j}]\big{\}}}{N(N-1)/2}\in[0,1].$ (12)
The three synthetic datasets used in this simulation are generated as in Li et al. (2022), and we report briefly their structures. For each $i=1,\ldots,N$, item $i$ has a true score $\gamma_{i}$, while the covariate vectors $\mathbf{x}_{i}=(x_{i1},\ldots,x_{iK_{x}})^{\prime}$ are drawn independently from a Normal distribution with zero mean and covariance equal to $\mathbb{C}ov(x_{il},x_{im})=\rho^{\left|l-m\right|}$, for $1\leq l,m\leq v$ and $\left|\rho\right|<1$. The true score vector $\bm{\gamma}$ is generated by considering the different roles of the covariates as:
  1. 1.
    $\gamma_{i}=\mathbf{x}_{i}^{\prime}\bm{\beta}$, where $\bm{\beta}=(3,2,-1,-0.5)^{\prime}$, $K_{x}=4$, and $\rho=0$;
  2. 2.
    $\gamma_{i}=\mathbf{x}_{i}^{\prime}\bm{\beta}+\left\lVert\mathbf{x}_{i}\right\rVert^{2}$, where $\bm{\beta}=(3,2,1)^{\prime}$, $K_{x}=3$, and $\rho=0.5$;
  3. 3.
    $\gamma_{i}=\left\lVert\mathbf{x}_{i}\right\rVert^{2}$, where $K_{x}=4$, and $\rho=0.5$.
The covariates are linearly related to the score in Scenario 1, have a nonlinear relationship in Scenario 3, and a combination of both in Scenario 2. Given the scores, $M$ full ranking lists $\{\tau_{j}\}_{j=1}^{M}$ are generated as $\tau_{j}=\operatorname{rank}(\mathbf{z}_{j})$ with $\mathbf{z}_{j}\mathbin{\overset{i.i.d.}{\kern 0.0pt\sim}}\mathcal{N}_{N}(\bm{\gamma},\sigma^{2}I_{N})$.
For each scenario and each value of $\sigma\in\{1,5,10,20,40\}$, we generate $100$ synthetic datasets with $N=50$ and $M=10$. We compare the results for ROBART against several competitors: the Borda Count, the Markov-Chain methods (MC1, MC2, MC3 - Dwork et al., 2001), the Cross-Entropy Monte Carlo method (CEMC - Lin and Ding, 2009), the Plankett-Luce (PL) model, the BARC model with and without covariates, and the BARC model with Mixture of rankers with different opinions (BARCM - Li et al., 2022). Further details on these methods are available in the Supplement.
We use the default values for the ROBART model as in Chipman et al. (2010) and follow Li et al. (2022) in choosing the hyperparameters for all the remaining models.
$\sigma$ Borda MC1 MC2 MC3 CEMC PL BAR $\text{BARC}_{1}$ $\text{BARC}_{2}$ $\text{BARC}_{3}$ BARCM ROBART Scenario 1 1 (0.03) 1.31 1.08 1.01 2.55 5.62 0.99 0.65 0.65 0.98 0.65 0.89 5 (0.13) 1.27 1.02 1.01 1.13 1.09 0.99 0.65 0.65 0.97 0.65 0.68 10 (0.22) 1.57 1.01 1.00 1.05 1.07 1.00 0.71 0.71 0.99 0.71 0.69 20 (0.32) 1.47 1.00 1.00 1.02 1.04 0.99 0.78 0.78 0.99 0.78 0.75 40 (0.41) 1.23 1.00 1.00 1.01 1.01 0.99 0.88 0.88 0.99 0.88 0.87 Scenario 2 1 (0.03) 1.33 1.04 1.00 2.65 1.06 0.99 0.97 0.97 0.98 0.97 0.86 5 (0.12) 1.30 1.01 1.00 1.12 1.10 0.99 0.83 0.83 0.97 0.83 0.69 10 (0.20) 1.36 1.01 1.00 1.05 1.07 1.00 0.77 0.77 0.99 0.77 0.67 20 (0.29) 1.48 1.01 1.00 1.03 1.04 0.99 0.80 0.80 0.99 0.80 0.72 40 (0.39) 1.25 1.00 1.00 1.01 1.03 1.00 0.89 0.89 1.00 0.89 0.85 Scenario 3 1 (0.05) 1.28 1.03 1.00 1.60 1.11 1.00 1.00 1.00 1.00 1.00 0.88 5 (0.18) 1.32 1.01 1.00 1.06 1.06 0.99 1.00 1.00 0.99 1.00 0.73 10 (0.29) 1.38 1.00 1.00 1.03 1.04 0.99 1.00 1.00 0.99 1.00 0.78 20 (0.36) 1.34 1.00 1.00 1.02 1.03 1.00 1.00 1.00 1.00 1.00 0.85 40 (0.42) 1.18 1.00 1.00 1.01 1.01 1.00 1.00 1.00 1.00 1.00 0.91  Notes: Scaled Kendall tau distance of rank predictions for the scenarios specified by Li et al. (2022). The Borda column with parentheses shows the average Kendall tau distances between estimated and true ranking lists for the Borda Count predictions, while the remaining columns show tau distances relative to the tau distance attained by the Borda Count method. Bold denotes the best-performing model.
Table 1: Comparison between ROBART, BARC, and other ranking methods. Full rankings.
Table 1 reports the Kendall tau distances between the true and estimated rank lists averaged over the $100$ replications. For each column, we provide the ratios of the average Kendall tau distances over the corresponding values for the Borda Count, which serves as the baseline. In Scenarios 2 and 3, for any $\sigma$, the proposed ROBART model outperforms all the competing models. In particular, there is a strong gain in Scenario 3, for any $\sigma$, moving from around 10% (for $\sigma=1$ or $40$) to 27% (for $\sigma=5$). Similarly in Scenario 2, irrespective of $\sigma$, ROBART performs better than BARC and BARCM, with a gain of 10% (for $\sigma$ equal to 1) to 20% (for $\sigma$ equal to 5). Increasing the value of $\sigma$ leads to similar conclusions, except for $\sigma$ equal to 40, where the ROBART model provides slightly better predictions than BARCM and BARC. These better performances are less evident in Scenario 1, where for small values of $\sigma$ (equal to 1 and 5), the ROBART is outperformed by the BARC and BARCM models, while it beats the other models of 1% and 3% when $\sigma$ is bigger than 10. As expected, nonlinear relationships in the data-generating process are correctly estimated by the BART specification. Unsurprisingly, the proposed model does not yield improvements for estimates of linear functions of covariates when $\sigma$ is small.
4.1.2 Static Simulation Study with Partial Rankings
ROBART and ARROBART are extendable to partial ranking data by replacing the constraints on the latent scores with the less restrictive constraints implied by partial rankings. The Supplement contains additional details for the model and sampler for partial ranking data. We demonstrate the ability of our model to predict rankings using partial ranking information through a simulation study described by Li et al. (2022). The comparison also included two additional specifications: the RODART, which is ROBART with a Dirichlet hyperprior on covariates’ splitting probabilities, and the SoftROBART, which is ROBART with soft and sparse splitting rules (Linero and Yang, 2018).
We divide the $N=80$ items into $K=1,2,4,8,10,16$ groups, of $N/K$ items. Pairwise comparison information is only available within groups. The DGP is otherwise the same as scenario 2 above. Table 2 contains the correlations of rank predictions averaged over 10 repetitions of the simulation study with latent error standard deviations $\sigma\in\{1,5\}$, as in Li et al. (2022). The results confirm that ROBART still provides a considerable improvement over linear models when only partial rankings are available. Finally, the SoftROBART is found to provide further improvements in ranking predictions.
K BAR BARC1 BARC2 BARC3 ROBART RODART SoftROBART $\sigma=1$ Scenario 2 1 (0.03) 1.00 1.00 1.00 0.83 0.84 0.67 2 (0.07) 0.61 0.60 0.59 0.32 0.34 0.28 4 (0.08) 0.59 0.62 0.64 0.35 0.39 0.28 8 (0.13) 0.51 0.49 0.56 0.24 0.27 0.21 10 (0.12) 0.59 0.59 0.65 0.26 0.33 0.22 16 (0.17) 0.55 0.56 0.58 0.23 0.36 0.21 $\sigma=5$ Scenario 2 1 (0.12) 0.83 0.83 0.98 0.60 0.68 0.46 2 (0.13) 0.80 0.79 1.04 0.54 0.64 0.46 4 (0.14) 0.78 0.78 0.95 0.56 0.73 0.46 8 (0.17) 0.63 0.64 0.84 0.46 0.58 0.36 10 (0.16) 0.65 0.65 0.88 0.45 0.57 0.37 16 (0.19) 0.60 0.60 0.78 0.42 0.58 0.35  Notes: Scaled Kendall tau distance of rank predictions for the scenarios specified by Li et al. (2022). The BAR column with parentheses shows the average Kendall tau distances between estimated and true ranking lists for the BAR predictions, while the remaining columns show tau distances relative to the tau distance attained by BAR.
RODART is ROBART with a Dirichlet hyperprior on covariates’ splitting probabilities. SoftROBART is ROBART with soft and sparse splitting rules.
Table 2: Comparison between ROBART and other ranking methods. Partial rankings.
4.1.3 Static Simulation Study with Ranker-specific Covariates
Table  3 contains the results for a variation on the static simulation study described by Li et al. (2022) in which the covariates are ranker-specific. For each $(i,m)\in\{1,\ldots,N\}\times\{1,\dots,M\}$, item $(i,m)$ has a true score $\gamma_{i,m}$, while the covariate vectors $\mathbf{x}_{i,m}=(x_{i,m,1},\ldots,x_{i,m,K_{x}})^{\prime}$ are drawn independently from a Normal distribution with zero mean and covariance equal to $\mathbb{C}ov(x_{i,m,l},x_{i,m,s})=\rho^{\left|l-s\right|}$, for $1\leq l,s\leq v$ and $\left|\rho\right|<1$. The true score vector $\bm{\gamma}$ is generated by considering the different roles of the covariates as:
  1. 1.
    $\gamma_{i,m}=\mathbf{x}_{i,m}^{\prime}\bm{\beta}$, where $\bm{\beta}=(3,2,-1,-0.5)^{\prime}$, $K_{x}=4$, and $\rho=0$;
  2. 2.
    $\gamma_{i,m}=\mathbf{x}_{i,m}^{\prime}\bm{\beta}+\left\lVert\mathbf{x}_{i,m}\right\rVert^{2}$, where $\bm{\beta}=(3,2,1)^{\prime}$, $K_{x}=3$, and $\rho=0.5$;
  3. 3.
    $\gamma_{i,m}=\left\lVert\mathbf{x}_{i,m}\right\rVert^{2}$, where $K_{x}=4$, and $\rho=0.5$.
ROBART provides much better predictions than other methods for all scenarios and values of $\sigma$. However, the benchmark methods were not originally intended for data with ranker-specific covariates.
$\sigma$ Borda MC1 MC2 MC3 CEMC PL BAR $\text{BARC}_{1}$ $\text{BARC}_{2}$ $\text{BARC}_{3}$ ROBART Scenario 1 1 (0.40) 1.27 1.00 1.00 1.00 1.03 1.27 1.27 1.27 1.27 0.09 5 (0.44) 1.15 1.00 1.00 1.01 1.02 1.15 1.15 1.15 1.15 0.20 10 (0.47) 1.08 1.00 1.00 1.00 1.00 1.08 1.08 1.08 1.08 0.31 20 (0.48) 1.03 1.00 1.00 1.00 1.00 1.03 1.03 1.03 1.03 0.51 40 (0.49) 1.02 1.00 1.00 1.00 1.00 1.02 1.02 1.02 1.02 0.72 Scenario 2 1 (0.40) 1.27 1.00 1.00 1.00 1.03 1.27 1.27 1.27 1.27 0.10 5 (0.43) 1.16 1.00 1.00 1.01 1.03 1.16 1.16 1.16 1.16 0.17 10 (0.46) 1.09 1.00 1.00 1.00 1.02 1.09 1.09 1.09 1.09 0.26 20 (0.48) 1.05 1.00 1.00 1.00 1.01 1.05 1.05 1.05 1.05 0.41 40 (0.49) 1.03 1.00 1.00 1.00 1.00 1.03 1.03 1.03 1.03 0.62 Scenario 3 1 (0.41) 1.25 1.00 1.00 1.01 1.04 1.25 1.25 1.25 1.25 0.13 5 (0.45) 1.10 1.00 1.00 1.01 1.02 1.10 1.10 1.10 1.10 0.27 10 (0.47) 1.05 1.00 1.00 1.00 1.01 1.05 1.05 1.05 1.05 0.44 20 (0.49) 1.03 1.00 1.00 1.00 1.00 1.03 1.03 1.03 1.03 0.64 40 (0.49) 1.02 1.00 1.00 1.00 1.00 1.02 1.02 1.02 1.02 0.79  Notes: Scaled Kendall tau distance of rank predictions for modified scenarios in which the covariates are ranker-specific. The Borda column with parentheses shows the average Kendall tau distances between estimated and true ranking lists for the Borda Count predictions, while the remaining columns show tau distances relative to the tau distance attained by the Borda Count method. Bold denotes the best-performing model.
Table 3: Comparison between ROBART, BARC, and other ranking methods. Full rankings with ranker-specific covariates.
4.2 Comparison of ARROBART and Linear Models
We investigate the performance of the proposed ARROBART against several competitors by means of the normalized Kendall tau distance defined in Eq. (12). In particular, the synthetic datasets are generated according to three main scenarios:
  1. 1.
    $\gamma_{ij,t}=0.1z_{ij,t-1}^{2}$,
  2. 2.
    $\gamma_{ij,t}=0.05z_{ij,t-1}+0.1z_{ij,t-1}^{2}$,
  3. 3.
    $\gamma_{ij,t}=0.1z_{ij,t-1}x_{ij,t-1,1}$,  $K_{x}=3$,  $\rho=0.5$,
where in the last scenario the $K_{x}$ covariates $\mathbf{x}_{ij,t}=(x_{ij,t,1},\ldots,x_{ij,t,K_{x}})^{\prime}$ are drawn independently from a multivariate normal distribution with mean zero and $\mathbb{C}ov(x_{ij,t,l},x_{ij,t,m})=\rho^{\left|l-m\right|}$, for $1\leq l,m\leq v$ and $\left|\rho\right|<1$. In detail, each scenario describes a nonlinear relationship, where $i=1,\ldots,N$, $j=1,\ldots,M$, and $t=1,\ldots,T$. We fix the number of rankers, $M$, the number of items, $N$ and the period, $T$, to $5$, $20$ and $52$, respectively.
For each scenario, the latent scores and the full ranking lists are generated as $\mathbf{z}_{j,t}\mathbin{\overset{iid}{\kern 0.0pt\sim}}\mathcal{N}_{N}(\bm{\gamma}_{jt},\sigma^{2}I_{N})$ and $\tau_{j,t}=\operatorname{rank}(\mathbf{z}_{j,t})$, respectively. Moreover, for each scenario, we generate three datasets varying the signal-to-noise ratio; specifically, the data are simulated using a value of $\sigma\in\{0.1,0.5,1.0,1.5,2.0\}$.
We compare the results for the full and the partial ranking against different competitors, such as (i) an autoregressive linear model for the latent score (ARROLinear); (ii) an ARROBART specification with exogenous covariates (ARROBARTX); (iii) an autoregressive linear model with exogenous covariates (ARROLinearX). Moreover, we extend the ARROBART model to include the lag-1 of the observed rank (ARROBART-lag) and compare it with (i) an analogous linear model (ARROLinear-lag); (ii) a ROBART model using the first lag of the response variable as a covariate (ROBART-lag); (iii) a linear rank order model (BARC) with the first lag of the response variable (ROLinear-lag).
Scenario 1 Scenario 2 Scenario 3 $\sigma$ 0.1 0.5 1.0 1.5 2.0 0.1 0.5 1.0 1.5 2.0 0.1 0.5 1.0 1.5 2.0 Full ranking ARROBART (0.38) (0.23) (0.22) (0.16) (0.15) (0.13) (0.18) (0.17) (0.15) (0.13) (0.54) (0.52) (0.49) (0.50) (0.52) ARROLinear 1.35 2.10 2.27 2.90 2.95 2.54 2.21 2.42 2.65 2.87 0.95 0.94 0.99 1.02 0.96 ARROBARTX 0.89 0.92 0.94 0.99 0.94 ARROLinearX 0.92 0.97 1.01 1.01 0.99 Partial ranking ARROBART (0.45) (0.36) (0.32) (0.29) (0.30) (0.19) (0.29) (0.26) (0.26) (0.27) (0.49) (0.49) (0.49) (0.50) (0.51) ARROLinear 1.13 1.31 1.55 1.71 1.65 2.06 1.62 1.83 1.90 1.70 1.01 1.02 1.01 0.96 1.00 ARROBARTX 0.97 1.02 1.00 0.95 0.95 ARROLinearX 1.00 0.99 1.03 0.97 0.99  Notes: The ARROBART row shows in brackets the average Kendall’s tau distances between estimated and true ranking lists using the ARROBART method. The remaining rows report the ratio of average Kendall’s tau for the model in the row to average Kendall’s tau for ARROBART. Values greater than one mean that the benchmark outperforms the competitor. Bold denotes the best-performing model.
Table 4: Comparison between ARROBART and ARROLinear with and without covariates for the full (top panel) and the partial (bottom panel) ranking.
Scenario 1 Scenario 2 Scenario 3 $\sigma$ 0.1 0.5 1.0 1.5 2.0 0.1 0.5 1.0 1.5 2.0 0.1 0.5 1.0 1.5 2.0 Full ranking ARROBART-lag (0.45) (0.34) (0.27) (0.19) (0.18) (0.19) (0.29) (0.25) (0.17) (0.15) (0.48) (0.42) (0.42) (0.41) (0.44) ARROLinear-lag 1.14 1.43 1.84 2.42 2.44 1.57 1.34 1.62 2.28 2.35 1.06 1.19 1.17 1.25 1.16 ROBART-lag 1.04 1.03 0.99 1.06 1.00 1.47 1.07 1.01 1.24 1.11 0.68 0.95 0.80 0.85 0.80 ROLinear-lag 1.14 1.45 1.82 2.40 2.43 1.53 1.33 1.62 2.29 2.37 1.06 1.21 1.17 1.25 1.16 With Irrelevant X ARROBART-lag 1.07 1.00 1.03 1.20 1.15 1.49 0.98 0.93 1.24 1.25 ARROLinear-lag 1.10 1.42 1.84 2.40 2.48 1.76 1.31 1.62 2.29 2.44 ROBART-lag 1.03 0.99 0.97 1.17 1.13 1.62 1.03 0.98 1.34 1.17 ROLinear-lag 1.10 1.44 1.84 2.38 2.49 1.73 1.29 1.62 2.27 2.41 Partial ranking ARROBART-lag (0.47) (0.42) (0.33) (0.32) (0.30) (0.30) (0.35) (0.30) (0.29) (0.28) (0.45) (0.47) (0.46) (0.48) (0.49) ARROLinear-lag 1.09 1.09 1.42 1.46 1.60 1.21 1.34 1.62 1.63 1.63 1.09 1.03 1.10 1.01 1.03 ROBART-lag 1.04 1.01 1.05 0.94 1.01 0.96 1.03 1.07 1.01 1.00 0.84 0.82 0.81 0.80 0.75 ROLinear-lag 1.08 1.08 1.44 1.47 1.64 1.22 1.37 1.65 1.68 1.68 1.10 1.03 1.11 1.01 1.04  Notes: The ARROBART-lag row shows in brackets the average Kendall’s tau distances between estimated and true ranking lists using the ARROBART-lag method. The remaining rows report the ratio of average Kendall’s tau for the model in the row to average Kendall’s tau for ARROBART-lag. Values greater than one mean that the benchmark outperforms the competitor. Bold denotes the best-performing model.
Table 5: Comparison between ARROBART, ARROLinear, ROBART, and ROLinear with one lag of the observed ranks as a covariate for the full (top panel) and the partial (bottom panel) ranking.
The results are shown in Table 4, where the normalized Kendall tau distance is reported for the benchmark ARROBART model, whereas for the competitors the ratio of their performance over the ARROBART is shown. Instead, Table 5 illustrates the outcome for all the models estimated using also the lag of the observed rank as a covariate, where we use the ARROBART-lag model as a benchmark. For the benchmark models, the closer the normalized Kendall tau distance is to zero, the better the performance. Therefore, ratios of Kendall’s tau distances for the competitors greater than 1 mean that the benchmark is outperforming them (and vice versa). The results suggest several interesting insights. First, the ARROBART (ARROBART-lag) benchmark outperforms the competitors for almost all the noise levels in scenarios 1 and 2. In scenarios 1 and 2, ARROBART-lag is almost uniformly better than all methods. The improvement over ARROLinear and Kendall’s tau for ARROBART are worse in the presence of a strong signal ($\sigma=0.1$); this is possibly due to the sub-optimal performance of the BART specification in almost the absence of noise and is in line with previous results in the literature (Chipman et al., 2010). Second, in the third scenario, the ARROBARTX (Tab. 4) and ROBART-lag (Tab. 5) are the best models. This demonstrates the ability of the proposed methods to leverage exogenous covariates when they are relevant to the regression. Third, the results suggest that the nonlinear specifications, that is the ARROBART and ROBART models, almost always outperform their respective linear counterparts. Finally, the results from Tab. 4 and  5 are confirmed even in the presence of partial ranking.
5 Real Data Applications
The usefulness of our static data models is demonstrated by an application to three standard label ranking benchmark datasets (de Sá et al., 2018). We include in these results a variation on ROBART, named SoftROBART, with soft and sparse splitting rules (Linero and Yang, 2018). Then, we compare static and dynamic model specifications on the two time series datasets presented in Section 1, about country-rankings on the economic complexity index and the 2022 NCAA rankings.
5.1 Static Label Rankings
The sushi dataset includes 5000 rankings of 10 items: shrimp, sea eel, tuna, squid, sea urchin, salmon roe, egg fatty tuna, tuna roll and cucumber roll (Kamishima et al., 2010). The variables include age, gender, answer duration, a binary variable indicating if the respondent moved to another city, and categorical variables for current and past region and prefecture. The categorical variables were re-coded as dummy variables resulting in 128 predictors.
The Top7Movies dataset is a subset of the MovieLens dataset (Harper and Konstan, 2016) processed by de Sá et al. (2018). The rankings were constructed from ratings of 7 movies submitted by 602 individuals. Predictors include gender, age, latitude and longitude of zipcode, city, state, and occupation. The categorical variables for city, state, and occupation were coded as dummy variables, resulting in a total of 446 predictors. This dataset contains many ties, and therefore allows for a comparison of methods in an application to partial ranking data.
The german2009 dataset was constructed from 2009 electoral results for 5 political parties across 413 districts. The predictors include an urban dummy variable, state, region, latitude, longitude, electoral participation rates, and a number of demographic and economic variables. The state and region variables were re-coded as dummy variables, resulting in 51 predictors in total. These rankings contain a small number of ties.
It can be observed in Table 6 below that ROBART and SoftROBART produce much more accurate ranking forecasts for the Top7Movies data than the benchmark linear models, and there is a smaller improvement for the sushi data. For the german2009 data, standard ROBART does not outperform linear models, whereas SoftROBART produces a comparable improvement to that achieved for the Top7Movies data. This suggests that there may be some smoothness or sparsity in the data generating process.
Table 6: Kendall’s tau distance between one-step ahead out-of-sample forecasted rankings and true rankings.
BARC1 BARC2 BARC3 ROBART SoftROBART
Sushi (0.34) 1.00 1.00 0.97 0.97
Top-7 Movies (0.44) 1.07 1.03 0.79 0.78
German Election 2009 (0.06) 0.94 0.95 1.09 0.76
  •  Notes: All methods were evaluated using 5-fold cross-validation. All results except the BARC1 results are relative to BARC1.
5.2 Dynamic Label Ranking: Economic Complexity Index
The ECI dataset (Alvo and Philip, 2014; Liu et al., 2019) includes data for 80 countries from 1965 to 2015. The dependent variable is the country-year specific ranking in an index defined by the Observatory of Economic Complexity. The independent variables include the lag of the observed rank, GDP per capita, population, population growth, cereal production in tonnes, the dependency ratio, and growth in the dependency ratio. These are not necessarily the most relevant variables for the index of interest. However, this exercise demonstrates the ability of the new methods to select relevant variables.
Table 7 shows that standard ROBART and ARROBART do not improve on a static linear model. However, RODART and ARRODART, variations on ROBART and ARROBART for which a Dirichlet hyperprior distribution is specified for sparse splitting rules (Linero, 2018), improve substantially on the linear models. Specifically, the dynamic ARRODART model provides better forecasts than the static RODART model. We refer to the Supplement for the data sources and further results.
Table 7: Kendall’s tau distance between one-step ahead out-of-sample forecasted rankings and true rankings.
ROLinear ARROLinear ROBART RODART ARROBART ARRODART
1.00 1.41 0.99 0.93 0.99 0.81
  •  Notes: All results except for the ROLinear result report the ratio of average Kendall’s tau for the model in the row to average Kendall’s tau for ROLinear. Values greater than one mean that ROLinear outperforms the competitor. Bold denotes the best-performing model.
5.3 Panel of Label Rankings: NCAA Pollsters
We apply the proposed methods to a real dataset of pollster-specific rankings of American football teams from the Associated Press weekly poll for the 2022 NCAA Division I Football season (see Fig. 1). We refer to the Supplement for the details of the data sources and further results. The dataset includes the ranking lists of $M=15$ pollsters across $T=16$ weeks. As many teams are not listed in the top 20 by at least one pollster in at least one week, we consider a subset of $N=7$ teams to obtain a time series of full rankings.
Figure 2: AP Pollster Brian Howell: observed poll ranks (blue dots and line) and ranks forecasted using the ARROBART (green triangles) and ARROLinear (orange squares) models.
To inspect in more detail the advantages of the BART specification compared to the linear one, we report the point forecasts from the best ARROBART and ARROLinear models in Fig. 2 against the observed ranks, for $N=7$ teams and $h=5$ test periods. As the rankers are assumed to be independent, we present the results for a single pollster. The ARROBART model is able to correctly predict the rank for several teams, with minor deviations for Alabama and Clemson. Conversely, the ARROLinear forecasts are generally more distant from the actual ranks. In particular, it can be observed that while both methods often produce similar forecasts, ARROBART produces more accurate forecasts for Alabama, USC, and Michigan in weeks 12 to 14.
The point forecast falls short of quantifying the uncertainty about the predicted value. To tackle this issue, we investigate in deeper detail the posterior predictive distribution of the best-performing model. Figure 3 reports the posterior predictive distribution of the rank for Alabama and Michigan, whose observed ranks have changed multiple times during the testing sample (see Fig. 2). For both teams, the predictive density of the rank is unimodal and changes over time in terms of its mode and dispersion, suggesting a horizon-specific uncertainty of predictions. Furthermore, the comparison of Fig. 2 and Fig. 3 highlights that in those periods where the point forecast differs from the observed one, the latter falls within the second-most frequent rank of the posterior predictive distribution.
Refer to caption Refer to caption
Figure 3: ARROBART posterior predictive distribution (grey balls, with size proportional to the probability mass) against the observed ranks (black dots and line) for Alabama (left) and Michigan (right) ranks of AP Pollster Brian Howell.
6 Discussion
This article introduces a novel class of nonparametric Thurstone models for static and dynamic regressions for rank-order data called ROBART and ARROBART. The proposed frameworks use a BART specification for the regression function of the latent variable representation to allow for flexible, possibly nonlinear relationships between the latent scores and their lagged values as well as exogenous covariates. The filtering and smoothing recursions for the dynamic model are derived, followed by a discussion on the computational tools needed to sample from them. A simulation study illustrates that the proposed ROBART and ARROBART outperform the competitors in forecasting. Then, ROBART is applied to rankings of sushi, movies, and political parties, and ARROBART is applied to a time series of economic complexity and NCAA rankings. The results highlight the better performance of the ROBART and ARROBART models compared to linear and static counterparts, suggesting the importance of accounting for temporal persistence and the usefulness of BART in accounting for the nonlinearities driving the evolution of the latent item-specific scores.
The proposed methods can be applied to a wide range of datasets. Examples include the analysis of the main drivers of customers’ preferences, university rankings, and political ratings, in addition to sports data. The proposed ARROBART model can be generalized to $p>1$ lags in the presence of more complex dynamics of the latent scores. Other interesting possible extensions of the proposed methods include investigating the performance of more general BART specifications. For instance, the introduction of a varying-coefficient BART (Deshpande et al., 2020) or partially linear BART model for the latent scores (Zeldow et al., 2019). The proposed autoregressive BART methods can also be adapted to investigate binary, ordered categorical, unordered categorical data, or censored outcome data. These models involve an autoregressive Gaussian process for the latent scores and are therefore similar to the rank-order model described in this article. The samplers can be derived in a similar way to the sampler in this work, and have simpler constraints on the latent variables than rank-order outcome models. Finally, it would be interesting to extend the proposed dynamic model to investigate the problem of aggregation of ranking lists (Deng et al., 2014; Zhu et al., 2023) in a time series setting.
Data Availability Statement
The data that support the findings of this study are openly available. These data were derived from the following resources: